### v1

# SpotLight: Shadow-Guided Object Relighting via Diffusion

Abstract
Recent work has shown that diffusion models can be used as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. Unlike typical physics-based renderers, however, neural rendering engines are limited by the lack of manual control over the lighting setup, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by specifying the desired shadows of the object. Rather surprisingly, we show that injecting only the shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. Specifically, we demonstrate its use with two neural renderers from the recent literature. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting.
| Unconditioned | Shadow-conditioned (top: shadow mapping, bottom: hand-drawn scribbles) composites from SpotLight | |||||
1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a single photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex surface interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [52]) has enabled exciting possibilities for image relighting [32, 73].
This paper focuses on providing local lighting control, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
In this paper, we present SpotLight, a method for realistic, controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than training a model specifically for the task of object relighting, our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow to a general-purpose neural renderer. Indeed, we find that simply specifying the desired lighting through the target shadow makes the renderer relight the object accordingly! SpotLight: Shadow-Guided Object Relighting via Diffusion shows two examples of such shadow-based lighting control: 1) the shadow of the couch is rendered using basic shadow mapping [19] according to the desired direction (top); and 2) one can even draw scribbles by hand to depict the desired shadow of the ottoman (bottom). To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained neural renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
The core technical contribution lies in our shadow injection approach which is robust to artifacts in the target shadow map. We build upon existing object compositing networks [75, 72] that operate on intrinsic images and exploit diffusion models’ learned priors to render realistic outputs. We demonstrate SpotLight handles complex scenarios and maintains physical plausibility without requiring extensive manual intervention nor any additional training.
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring fine-tuning or additional training. We demonstrate superior quality, both perceptually through a user study and quantitatively, compared to competitive specialized models. Finally, we release a novel evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [58, 28, 59, 52]. These models offer control mechanisms for various modalities, such as text [48, 52], image content [53, 21], intrinsic images [75, 42, 72], classifier guidance [16] and cross-modal data [12, 69]. Notably, training-free guidance methods [43, 20, 3, 4] have enabled image editing capabilities without specific training, forming the basis of our approach.
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for style transfer [25, 41] or harmonization [51, 46, 35, 61] do not prioritize physically accurate shading of the objects. Recently, [47] utilized a single-view multi-illumination dataset [44] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [65, 11] for relighting indoor scenes. IC-Light [73] enforces consistency in appearance to relight portraits and various objects. Some methods specialize in outdoor scenes, using geometric priors like depth [27] or normalized coordinates [32] or both [38] to improve global lighting estimation. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. ZeroComp [75] learns to composite virtual objects in a zero-shot fashion, but does not offer lighting control. RGBX [72] uses diffusion models for material- and lighting-aware composition with text-guided control. Neural Gaffer [30] and DiLightNet [71] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow guidance for object relighting.
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 37, 76], though they generally lack controllability. Parametric models offer more control [23, 63, 15], but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
Shadow generation Generative models for rendering realistic shadows [74, 39, 29, 40], or harmonizing rendered shadows with the image [62] have been proposed. Other methods enable controllable shadows for 3D [55] and 2D objects [56, 57]. Our approach can leverage existing shadow generation methods and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 34, 45, 64, 13]. Many works have also focused on estimating scene geometry, for example, depth [49, 8, 70, 31, 66, 67] and normals [18, 6, 50, 5, 68]. Recently, approaches such as [37, 76] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 31, 68, 2, 42]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
3 SpotLight
Diffusion-based neural renderers are currently limited by their inability to directly control shading, especially for specific objects as they are generally conditioned on albedo, normals, depth, and global shading. SpotLight addresses this limitation by enabling control over a single object lighting effects by defining a coarse shadow. Leveraging priors from pre-trained models and the diffusion framework, we introduce a method capable of harmonizing the object realistically from a provided shadow region. Fig. 2 presents an overview of the method, with each component described in detail below.
3.1 Rendering from intrinsics
Our method leverages a pre-trained diffusion-based neural renderer (e.g., ZeroComp [75], RGBX [72]) which accepts a multi-channel input of intrinsic maps —typically: albedo, shading, surface normals and depth, although other maps such as roughness and metallic can also be used—and outputs an RGB image consistent with these inputs. Our approach also requires shadow and object masks, and resp., where 1 indicates if the pixel is in shadow/object, and 0 otherwise. The discussion below assumes is based on latent diffusion [52], which performs the diffusion process in a low resolution latent space defined by a VAE encoder .
3.2 Blending shadows with the background
Provided a user defined shadow mask (), we first need to ensure that the diffusion model will properly blend the shadow in its final render. To achieve a perceptually pleasing shadow, we take inspiration from Blended Latent Diffusion [4] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadows composited over the image, . In practice, we obtain by compositing the object albedo and the shadow on the background. The update made at each timestep to incorporate the desired shadow latents is therefore
where is the shadow latent weight and is the shadow mask, first dilated by a square kernel to maintain the desired shadow softness on the edges, then bilinearly downsampled to the latent resolution (). Noise is added to the encoded composite using the DDIM [59] noise scheduler. Different from [4], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
3.3 Enhancing lighting control
Simply running a neural renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec. 4.6). We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by . Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample. We obtain the outputs of the diffusion renderer, which are typically -predictions [54], that is, a linear combination of the noise and the estimated image. We then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After applying eq. 2, we run a regular DDIM sampling step to obtain the latents .
3.4 Diffusion-based neural renderer backbone
Since SpotLight only applies changes to the latent space of a diffusion model, it can readily be applied to pre-trained diffusion-based neural renderers without any finetuning.
ZeroComp [75] provides a maskable shading map as input to a ControlNet-based neural renderer. Here, we use their approach directly and mask out the shading on both the object and the shadow region.
RGBX [72] provides a checkpoint of their XRGB model finetuned for inpainting rectangular masks. We mask out the bounding box including the object and shadow region.
At each denoising step, the denoising network is ran to obtain
where represents the network -prediction, represents the composited intrinsics. For ZeroComp, those intrinsics correspond to albedo, normals, depth, and partial shading and for RGBX, we use the albedo, normals, metallic, roughness and masked image.
3.5 Obtaining the final render
After having performed the denosing steps, we obtain an image by decoding the latents into by using the VAE decoder. The final result is obtained with the background preservation strategy proposed in [75], which obtains a shadow matte by computing the ratio of the image obtained with and without shadow guidance. This ensures the diffusion process does not affect the background image.
4 Evaluation on 3D object compositing
In this section, we quantitatively evaluate SpotLight against other methods that support lighting control on composition of 3D objects in 2D images. Additionally, we conduct a user study to assess human perception, comparing its performance with these competing approaches.
4.1 Evaluation dataset
We start from the evaluation dataset proposed in [75], where high-quality 3D models from the Amazon Berkeley Objects dataset [14] are inserted into background images cropped with a 50° field of view from the Laval Indoor HDR dataset [22]. The authors of [75] graciously provided us with reproducible Blender scene information, which we extend in two ways. First, we improve the shadow catcher using back-projected background depth estimated using Depth Anything V2 [67]. Second, object intrinsics are rendered with anti-aliased alpha mattes, allowing for more realistic compositing results. We use a depth-warping strategy on the HDR panorama as in [75] to account for spatially varying lighting around the object position. The dataset is rendered using the physically-based Cycles renderer with 2048 samples per pixel, a noise threshold of 0.01, and denoised using OpenImageDenoise [1]. We produce two versions of the dataset, each intended for different purposes.
“Reference-based” dataset This version identifies the dominant light source in the scene, extracted from the ground-truth warped panorama using the light extraction algorithm from [23]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3. We filter out samples where the background depth estimation failed. This data set is intended for quantitative evaluation (sec. 4.4), allowing the simulated ground truth of the inserted object to be used to compute metrics.
“User-controlled” dataset This second dataset simulates a scenario where a user specifies a desired light position. In this setup, five lighting directions are defined from left to right in increments in azimuth, with the light source passing behind the object to ensure the shadow remains visible. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,050 images. This dataset will be used in the user study (sec. 4.5) to assess the perceptual accuracy of renders in edited lighting scenarios.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) | Simulated GT |
4.2 SpotLight specifics
As mentioned in sec. 3.4, our proposed SpotLight readily applies to existing diffusion-based neural renderers. Here, we experiment with two approaches from the recent literature: ZeroComp [75] and RGBX [72].
ZeroComp backbone We train ZeroComp following the procedure outlined in [75], and use the same conditioned intrinsic estimators at inference time: StableNormal [68] for normals, IID [33] for albedo, and ZoeDepth [8] for depth. The masked shading in the background is obtained by dividing the image by the albedo and applying the same masking strategy. Given its superior results, we adopt this method as our default model for 3D object compositing.
RGBX backbone We use the RGBX model to extract intrinsic properties (normals, metallic, roughness, and albedo). We observe that results are generally overly bright, likely due to a domain gap between the training and evaluation images. To address this, we re-expose the background by a factor of 2 before feeding it to the network, then divide the output by 2, in linear space. We apply the same background preservation strategy as in ZeroComp.
Light control SpotLight requires a coarse shadow for guidance. Therefore, we use a real-time rasterization renderer (Blender’s EEVEE) to compute a simple shadow of the 3D object. This shadow is generated from a simple parametric distant sphere light and projected onto the mesh constructed from the estimated depth map of the background.
4.3 Baselines
We compare SpotLight against recent diffusion-based methods that support light control: DiLightNet [71], IC-Light [73], and Neural Gaffer [30], using their public implementations. While these baselines can relight objects, none are capable of generating a shadow around the object. For fairness, the physically-based renderer Cycles is therefore used to cast a realistic shadow. OpenImageDenoise [1] is also used to denoise the results of all methods. We begin by describing how light control is given to baselines, then proceed to describe per-method specifics.
4.3.1 Light control
In order to ensure fairness in evaluation, we aim to provide each method with a light parametrization that is as close as possible to the ground truth dominant light direction available in the evaluation dataset.
Both DiLightNet [71] and Neural Gaffer [30] require an environment map for lighting. We construct such an environment map with radiance along direction defined as the sum of a spherical gaussian and a constant term
where and are the RGB colors of the light and ambient terms resp., is the dominant light source direction, and is the bandwidth. is obtained by computing the average color over the background image. Panoramas from the Laval Indoor HDR dataset (excluding those in our test set) are used to estimate a single average intensity of the dominant light source, , which is set to the ratio of the integral of the brightest pixels in the panoramas, divided by the integral over all pixels. We further divide this integral by 2 to consider a single hemisphere and avoid overly bright dominant light sources. The light color is defined by , where is the normalized ambient color . We found that the bandwidth parameter did not advantage any specific method and therefore fixed it to , which we typically observe for an indoor light.
4.3.2 Per-method details
DiLightNet [71] DiLightNet requires a prompt describing the object, which we obtain by feeding the ground truth object composited on a black background to BLIP-2 [36] with the prompt “Putting aside the black background, this object is a”. To generate the results, depth and radiance hints are required. We supply DiLightNet with depth information from the 3D model of the object and render the radiance hints using the input environment map (sec. 4.3.1) as outlined in [71]. The rendered object is then composited with the background, along with the rendered shadow.
Neural Gaffer [30] We provide the method with the environment map described in sec. 4.3.1. The ground truth object is composited on a white background before feeding it to the network, as expected by the method. Since Neural Gaffer is trained at lower resolution (), its result is upsampled to then composited on the full resolution background. An erosion using a kernel is applied to the ground truth mask to mitigate boundary artifacts.
IC-Light [73] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning. We provide a rendering of the object lit by the ground truth environment map, giving a significant advantage to the method. No prompts are used.
4.4 Experimental results
Tab. 1 presents quantitative results comparing the various baselines (c.f. sec. 4.3) with two versions of SpotLight (c.f. sec. 4.2) on our “reference-based” dataset, where simulated ground truth is available (sec. 4.1). Here, all methods are conditioned on the dominant light direction using their specific light parametrization (sec. 4.3.1). It is worth noting that all three baselines (DiLightNet, IC-Light and Neural Gaffer) are given as input the ground truth object foreground and shadowed background, giving them a significant advantage. Fig. 3 shows qualitative results for each method. We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| DiLightNet [71]∗ | 25.36 | 0.950 | 0.059 | 0.024 | 0.044 |
| IC-Light [73]∗ | 28.59 | 0.965 | 0.043 | 0.018 | 0.034 |
| Neural Gaffer [30]∗ | 29.64 | 0.963 | 0.037 | 0.016 | 0.038 |
| SpotLight (RGBX) | 26.44 | 0.956 | 0.052 | 0.018 | 0.041 |
| SpotLight (ZeroComp) | 30.68 | 0.973 | 0.033 | 0.012 | 0.031 |
4.5 User study
A user study was conducted to evaluate the realism of local lighting control. Here, we use the “user-controlled” version of the dataset (see sec. 4.1) to evaluate whether the realism of results obtained by conditioning on a specific lighting direction, even if it is not the dominant light direction.
We evaluate the three methods with the best PSNR from tab. 1, namely SpotLight (with the ZeroComp backbone), Neural Gaffer [30] and IC-Light [73]. Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side. The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 40 random images are chosen, resulting in 120 pairs of images observed by each participant. We randomly shuffle the 120 pairs for each user, as well as the left-right order. Examples are shown in fig. 5.
The user study was completed by observers. Following Giroux et al. [26], we used the Thurstone Case V Law of Comparative Judgement [60] to obtain a scaling -score for each method, where a higher scale indicates a higher human preference. The results, shown in fig. 4 show that our method is the favorite. The 95% confidence interval error bars also show that the margin in performance is statistically significant, indicating that SpotLight achieves the highest realism.
| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) |
4.6 Effect of parameter selection
Our method employs parameters that control the degree of shadow guidance ( in eq. 2) and the shadow blending ( in eq. 1). Tab. 2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the simulated ground truth is available (c.f. sec. 4.1). Although using no guidance () results in better quantitative performance, we observe qualitatively in fig. 6 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control. Employing no shadow blending () results in similar metric values but makes shadows much less visible (fig. 7).
| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
| , (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.031 |
| (no guidance) | 31.65 | 0.976 | 0.030 | 0.011 | 0.029 |
| 28.60 | 0.966 | 0.044 | 0.015 | 0.036 | |
| 30.78 | 0.974 | 0.033 | 0.012 | 0.030 | |
| 29.32 | 0.970 | 0.039 | 0.014 | 0.034 |
| No guidance () | (ours) |
| (no blending) | (ours) |
5 Extensions
This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
Shadow control on 2D object composition We demonstrate that SpotLight can be applied to 2D images by leveraging a ZeroComp [75] backbone trained without depth maps, using the normals [68] and albedo [33] estimators. Without access to a 3D model, we cannot rely on a rendering engine to cast shadows. Instead, we employ PixHt-Lab [57] to generate realistic soft shadows from a controllable point light coordinate. Here, a sample without shadow is provided for negative guidance , see sec. 3.3. To estimate the required pixel height of the object cutout, we attribute each pixel -coordinate as the pixel height for demonstration purposes, more sophisticated methods such as [56] could be used. After rendering with the backbone, we enhance the target by adding the high-frequency details from the source image. This is achieved by subtracting a Gaussian-blurred version of the source image ( = 1) from itself (as in unsharp masking) and adding the resulting details to the target. Example 2D compositing results are shown in fig. 8, which demonstrate realistic 2D object relighting from a user-controlled shadow.
| Object | Background | Light-controllable composites | ||
Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and pixel height in 2D. However, one can simply draw the desired shadow! Fig. 9 and SpotLight: Shadow-Guided Object Relighting via Diffusion (first row) show that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
6 Discussion
We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training. Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
While the model can realistically estimate relighting from a shadow, generating a shadow map from scratch remains challenging. A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights. Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
Acknowledgements
This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix. Computing resources were provided by the Digital Research Alliance of Canada. The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.

### v2

@@
 Abstract
-Recent work has shown that diffusion models can be used as powerful neural rendering engines that can be leveraged for inserting virtual objects into images.
-Unlike typical physics-based renderers, however, neural rendering engines are limited by the lack of manual control over the lighting setup, which is often essential for improving or personalizing the desired image outcome.
-In this paper, we show that precise lighting control can be achieved for object relighting simply by specifying the desired shadows of the object.
-Rather surprisingly, we show that injecting only the shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image.
+Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images.
+However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome.
+In this paper, we show that precise lighting control can be achieved for object relighting simply by providing a coarse shadow of the object.
+Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image.
 Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training.
-Specifically, we demonstrate its use with two neural renderers from the recent literature.
 We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting.
-| Unconditioned | Shadow-conditioned (top: shadow mapping, bottom: hand-drawn scribbles) composites from SpotLight | |||||
+We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
+Background image | Previous method | SpotLight enables shadow-guided lighting control |
 1 Introduction
 Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process.
-When one is given a single photograph, however, no such control is possible.
-Relighting then becomes a hard task because of the complex surface interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture.
+When one is given a photograph, however, no such control is possible.
+Relighting then becomes a hard task because of the complex interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture.
 The recent rise of powerful image generation models (e.g., [54]) has enabled exciting possibilities for image relighting [31, 76].
-This paper focuses on providing local lighting control, where one wishes to modify the appearance of a single object.
+This paper focuses on providing lighting control to recent diffusion image relighting frameworks, where one wishes to modify the appearance of a single object.
 This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image.
 In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
-In this paper, we present SpotLight, a method for realistic, controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer.
-Rather than training a model specifically for the task of object relighting, our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow to a general-purpose neural renderer.
-Indeed, we find that simply specifying the desired lighting through the target shadow makes the renderer relight the object accordingly!
-SpotLight: Shadow-Guided Object Relighting via Diffusion shows two examples of such shadow-based lighting control: 1) the shadow of the couch is rendered using basic shadow mapping [19] according to the desired direction (top); and 2) one can even draw scribbles by hand to depict the desired shadow of the ottoman (bottom).
-To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained neural renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
-The core technical contribution lies in our shadow injection approach which is robust to artifacts in the target shadow map.
-We build upon existing object compositing networks [75, 72] that operate on intrinsic images and exploit diffusion models’ learned priors to render realistic outputs.
-We demonstrate SpotLight handles complex scenarios and maintains physical plausibility without requiring extensive manual intervention nor any additional training.
+We present SpotLight, a method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer.
+Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [74, 29]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer.
+Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it to relight the object accordingly.
+We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object as it moves around it.
+Fig.
+1 shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [19] according to the desired lighting direction.
+From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting.
+To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
+Our core technical contribution lies in our shadow guidance approach, which refines the guiding shadow input while realistically relighting the virtual object.
+Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control.
+This builds upon existing object compositing networks [78, 75] that operate on intrinsic images along with diffusion models’ learned priors to render realistic outputs.
+We demonstrate that SpotLight’s guidance mechanism handles complex scenarios and maintains physical plausibility without requiring manual intervention nor any additional training.
 In summary, our contributions are as follows.
-First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring fine-tuning or additional training.
-We demonstrate superior quality, both perceptually through a user study and quantitatively, compared to competitive specialized models.
-Finally, we release a novel evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
+First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring additional training.
+We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models.
+Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
 2 Related work
 Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [59, 27, 60, 54].
-These models offer control mechanisms for various modalities, such as text [48, 52], image content [53, 21], intrinsic images [75, 42, 72], classifier guidance [16] and cross-modal data [12, 69].
-Notably, training-free guidance methods [43, 20, 3, 4] have enabled image editing capabilities without specific training, forming the basis of our approach.
+These models offer control mechanisms for various modalities, such as text [50, 54], image content [55, 21], intrinsic images [78, 43, 75, 39], classifier guidance [16] and cross-modal data [12, 72].
+Notably, training-free guidance methods [44, 20, 3, 4, 26] have enabled image editing capabilities without specific training, forming the basis of our approach.
 Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials.
-Previous methods for style transfer [25, 41] or harmonization [51, 46, 35, 61] do not prioritize physically accurate shading of the objects.
+Previous methods for harmonization [53, 48, 34, 62] do not prioritize physically accurate shading of the objects.
 Recently, [49] utilized a single-view multi-illumination dataset [45] to enable direct control over the dominant lighting direction.
@@
 IC-Light [76] enforces consistency in appearance to relight portraits and various objects.
-Some methods specialize in outdoor scenes, using geometric priors like depth [27] or normalized coordinates [32] or both [38] to improve global lighting estimation.
+LumiNet [68] and ScribbleLight [14] achieve image relighting through illumination transfer or user-defined scribbles, respectively.
+Some methods specialize in outdoor scenes, using geometric priors like depth [25] or normalized coordinates [31] or both [40].
 These methods however cannot handle lighting on specific region or objects in the scene.
 Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13].
-ZeroComp [75] learns to composite virtual objects in a zero-shot fashion, but does not offer lighting control.
-RGBX [72] uses diffusion models for material- and lighting-aware composition with text-guided control.
-Neural Gaffer [30] and DiLightNet [71] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images.
-We bridge this gap by introducing a shadow guidance for object relighting.
-Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 37, 76], though they generally lack controllability.
-Parametric models offer more control [23, 63, 15], but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
+Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [78, 75, 39], which can be leveraged to achieve zero-shot object composition.
+However, these methods have limited lighting control.
+ZeroComp [78] provides no lighting control, RGBX [75] can only provide lighting control through a text prompt, and DiffusionRenderer [39] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set.
+Other diffusion-based approaches, such as Neural Gaffer [29], DiLightNet [74] and IllumiNeRF [79] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images.
+We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
+Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 36, 80, 47, 38], though they generally lack controllability.
+Parametric models [23, 65, 15] or GAN inversion methods [64] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
 Shadow generation Generative models for rendering realistic shadows [77, 41, 28, 42], or harmonizing rendered shadows with the image [63] have been proposed.
 Other methods enable controllable shadows for 3D [56] and 2D objects [57, 58].
-Our approach can leverage existing shadow generation methods and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
+Our approach can leverage existing shadow generation methods (e.g., [58], see supp.) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
 Image intrinsics Decomposing images into albedo and shading has long been studied [7, 33, 46, 66, 13].
-Many works have also focused on estimating scene geometry, for example, depth [49, 8, 70, 31, 66, 67] and normals [18, 6, 50, 5, 68].
+Numerous works have also focused on estimating scene geometry, for example, depth [51, 8, 73, 30, 69, 70] and normals [18, 6, 52, 5, 71].
 Recently, approaches such as [36, 80] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials.
@@
 We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
+3 Background: diffusion renderers
+Recent work [75, 78] has shown that diffusion models can be adapted to generate images from intrinsic maps including materials (e.g., albedo, roughness, metallic), geometry (e.g., surface normals and depth), and shading.
+This capability was built using a ControlNet with ZeroComp [78] or full finetuning with RGBX [75].
+We refer to this new class of conditional generation models as “diffusion renderers”.
+In these models, the denoising network is ran at each denoising step to obtain
+where represents the network -prediction, the noisy latents, and the conditioning intrinsic maps.
+It has been shown that these methods can realistically composite virtual objects into images in a zero-shot fashion, despite not being trained for this task.
+Provided a mask of the virtual object and its intrinsic maps (rendered with graphics shaders), the object can be trivially composited in most intrinsic maps of the background image (estimated with trained estimators).
+Since computing the virtual object shading requires complex lighting simulations, diffusion approaches estimate the shading by providing a partial shading map [78], or train inpainting models [75].
+However, unlike traditional rendering, these methods offer no control over the lighting conditions.
+In this work, we use a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, analogous to a spotlight moving around the object.
+We now describe our approach in detail.
 4 SpotLight
-Diffusion-based neural renderers are currently limited by their inability to directly control shading, especially for specific objects as they are generally conditioned on albedo, normals, depth, and global shading.
-SpotLight addresses this limitation by enabling control over a single object lighting effects by defining a coarse shadow.
-Leveraging priors from pre-trained models and the diffusion framework, we introduce a method capable of harmonizing the object realistically from a provided shadow region.
+We employ a user-provided, guiding shadow to steer the pre-trained diffusion process, thereby enforcing matching illumination on the inserted object.
+Our key insight is that even an approximate shadow encapsulates important lighting information.
+By integrating this cue early in the diffusion process, we simultaneously refine both the composite appearance and its shadow, yielding a natural blend with the scene.
+4.1 Approach overview
+Our approach requires two inputs: an object mask and a user-specified guiding shadow .
+This shadow, computed via fast rasterization or specified through hand-drawn scribbles, is used to guide the latent diffusion process.
+Without retraining the pre-trained renderer , we steer its generative process by incorporating the guiding shadow into the latent space.
+4.2 Blending shadows
+To fuse the guiding shadow with the latent representation, we update the noisy latent code at each diffusion timestep, similar to [4]. In our framework, the updated latent
+serves as a soft constraint that steers the process toward a shadow-consistent output.
+Here, is the composite of the object albedo, the target shadow, and the background, and controls the strength of the guiding shadow.
+The latent shadow mask is obtained by downsampling the target shadow to the latent resolution (see supp.
+for details).
+4.3 Shading the object via dual-branch guidance
+A key element of our approach is to amplify the impact of the guiding shadow on the object shading through classifier-free guidance (CFG) [26].
+We process two parallel diffusion branches: a positive branch that incorporates the desired shadow direction, and a negative branch that suppresses an undesired shadow direction (e.g., via an opposing light direction).
 Fig.
-2 presents an overview of the method, with each component described in detail below.
-3.1 Rendering from intrinsics
-Our method leverages a pre-trained diffusion-based neural renderer (e.g., ZeroComp [75], RGBX [72]) which accepts a multi-channel input of intrinsic maps —typically: albedo, shading, surface normals and depth, although other maps such as roughness and metallic can also be used—and outputs an RGB image consistent with these inputs.
-Our approach also requires shadow and object masks, and resp., where 1 indicates if the pixel is in shadow/object, and 0 otherwise.
-The discussion below assumes is based on latent diffusion [52], which performs the diffusion process in a low resolution latent space defined by a VAE encoder .
-3.2 Blending shadows with the background
-Provided a user defined shadow mask (), we first need to ensure that the diffusion model will properly blend the shadow in its final render.
-To achieve a perceptually pleasing shadow, we take inspiration from Blended Latent Diffusion [4] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadows composited over the image, .
-In practice, we obtain by compositing the object albedo and the shadow on the background.
-The update made at each timestep to incorporate the desired shadow latents is therefore
-where is the shadow latent weight and is the shadow mask, first dilated by a square kernel to maintain the desired shadow softness on the edges, then bilinearly downsampled to the latent resolution ().
-Noise is added to the encoded composite using the DDIM [59] noise scheduler.
-Different from [4], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
-3.3 Enhancing lighting control
-Simply running a neural renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec.
-4.6).
-We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by .
-Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample.
-We obtain the outputs of the diffusion renderer, which are typically -predictions [54], that is, a linear combination of the noise and the estimated image.
-We then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
-where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After applying eq. 2, we run a regular DDIM sampling step to obtain the latents .
-3.4 Diffusion-based neural renderer backbone
-Since SpotLight only applies changes to the latent space of a diffusion model, it can readily be applied to pre-trained diffusion-based neural renderers without any finetuning.
-ZeroComp [75] provides a maskable shading map as input to a ControlNet-based neural renderer. Here, we use their approach directly and mask out the shading on both the object and the shadow region.
-RGBX [72] provides a checkpoint of their XRGB model finetuned for inpainting rectangular masks. We mask out the bounding box including the object and shadow region.
-At each denoising step, the denoising network is ran to obtain
-where represents the network -prediction, represents the composited intrinsics.
-For ZeroComp, those intrinsics correspond to albedo, normals, depth, and partial shading and for RGBX, we use the albedo, normals, metallic, roughness and masked image.
-3.5 Obtaining the final render
-After having performed the denosing steps, we obtain an image by decoding the latents into by using the VAE decoder.
-The final result is obtained with the background preservation strategy proposed in [75], which obtains a shadow matte by computing the ratio of the image obtained with and without shadow guidance.
-This ensures the diffusion process does not affect the background image.
+2 shows our dual-branch guidance mechanism.
+Classifier-free guidance then combines these branches:
+where sets the guidance scale and is the object mask downsampled to the latent resolution.
+This dual-branch strategy reinforces the intended lighting on the object while naturally blending it with the background.
+4.4 Final image synthesis
+After the guided diffusion process, the resulting latent is decoded with the VAE decoder.
+The background preservation strategy of [78] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
 5 Evaluation on 3D object compositing
-In this section, we quantitatively evaluate SpotLight against other methods that support lighting control on composition of 3D objects in 2D images.
-Additionally, we conduct a user study to assess human perception, comparing its performance with these competing approaches.
+Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene.
+To address this, we build over the evaluation setup of [78] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas.
+We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render.
+In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
 5.1 Evaluation dataset
-We start from the evaluation dataset proposed in [75], where high-quality 3D models from the Amazon Berkeley Objects dataset [14] are inserted into background images cropped with a 50° field of view from the Laval Indoor HDR dataset [22].
-The authors of [75] graciously provided us with reproducible Blender scene information, which we extend in two ways.
-First, we improve the shadow catcher using back-projected background depth estimated using Depth Anything V2 [67].
-Second, object intrinsics are rendered with anti-aliased alpha mattes, allowing for more realistic compositing results.
-We use a depth-warping strategy on the HDR panorama as in [75] to account for spatially varying lighting around the object position.
-The dataset is rendered using the physically-based Cycles renderer with 2048 samples per pixel, a noise threshold of 0.01, and denoised using OpenImageDenoise [1].
-We produce two versions of the dataset, each intended for different purposes.
-“Reference-based” dataset This version identifies the dominant light source in the scene, extracted from the ground-truth warped panorama using the light extraction algorithm from [23].
+We create datasets for the two scenarios using the Blender scene information graciously provided by the authors of [78].
+We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [70].
+We also improve the ground truth intrinsic maps by providing correct antialiased versions.
+“Reference-based” scenario To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [78] to ensure lighting fidelity around the inserted object.
+We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [19].
 A total of 210 images featuring various objects are rendered, with example samples shown in fig.
-3.
-We filter out samples where the background depth estimation failed.
+3, in the “ground truth” column.
 This dataset is intended for quantitative evaluation (sec.
-4.4), allowing the simulated ground truth of the inserted object to be used to compute metrics.
-“User-controlled” dataset This second dataset simulates a scenario where a user specifies a desired light position.
-In this setup, five lighting directions are defined from left to right in increments in azimuth, with the light source passing behind the object to ensure the shadow remains visible.
-The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,050 images.
-This dataset will be used in the user study (sec.
-4.5) to assess the perceptual accuracy of renders in edited lighting scenarios.
-| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) | Simulated GT |
+5.4).
+“User-controlled” scenario This dataset simulates a scenario where a user specifies a desired light position.
+8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine.
+The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images.
+This dataset is utilized in the user study (sec.
+5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.
 5.2 SpotLight specifics
 As mentioned in sec.
-3.4, our proposed SpotLight readily applies to existing diffusion-based neural renderers.
-Here, we experiment with two approaches from the recent literature: ZeroComp [75] and RGBX [72].
-ZeroComp backbone We train ZeroComp following the procedure outlined in [75], and use the same conditioned intrinsic estimators at inference time: StableNormal [68] for normals, IID [33] for albedo, and ZoeDepth [8] for depth.
-The masked shading in the background is obtained by dividing the image by the albedo and applying the same masking strategy.
+3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [78] or RGBX [75].
+Here, we provide details on ZeroComp, see supp.
+for RGBX.
+We use the checkpoint trained on the Openrooms dataset [37] provided by the authors.
+The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions.
 Given its superior results, we adopt this method as our default model for 3D object compositing.
-RGBX backbone We use the RGBX model to extract intrinsic properties (normals, metallic, roughness, and albedo).
-We observe that results are generally overly bright, likely due to a domain gap between the training and evaluation images.
-To address this, we re-expose the background by a factor of 2 before feeding it to the network, then divide the output by 2, in linear space.
-We apply the same background preservation strategy as in ZeroComp.
-Light control SpotLight requires a coarse shadow for guidance.
-Therefore, we use a real-time rasterization renderer (Blender’s EEVEE) to compute a simple shadow of the 3D object.
-This shadow is generated from a simple parametric distant sphere light and projected onto the mesh constructed from the estimated depth map of the background.
+To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [78], namely StableNormal [71] for normals, IID [32] for albedo, and ZoeDepth [8] for depth.
+Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
+Light-conditioned methods | |||||
+DiLightNet [74] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
+Neural Gaffer [29] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
+Shadow-conditioned methods | |||||
+IC-Light [76] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
+ZeroComp+SDEdit [44] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
+SpotLight | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
 5.3 Baselines
-We compare SpotLight against recent diffusion-based methods that support light control: DiLightNet [71], IC-Light [73], and Neural Gaffer [30], using their public implementations.
-While these baselines can relight objects, none are capable of generating a shadow around the object.
-For fairness, the physically-based renderer Cycles is therefore used to cast a realistic shadow.
-OpenImageDenoise [1] is also used to denoise the results of all methods.
-We begin by describing how light control is given to baselines, then proceed to describe per-method specifics.
-4.3.1 Light control
-In order to ensure fairness in evaluation, we aim to provide each method with a light parametrization that is as close as possible to the ground truth dominant light direction available in the evaluation dataset.
-Both DiLightNet [71] and Neural Gaffer [30] require an environment map for lighting.
-We construct such an environment map with radiance along direction defined as the sum of a spherical gaussian and a constant term
-where and are the RGB colors of the light and ambient terms resp., is the dominant light source direction, and is the bandwidth.
-is obtained by computing the average color over the background image.
-Panoramas from the Laval Indoor HDR dataset (excluding those in our test set) are used to estimate a single average intensity of the dominant light source, , which is set to the ratio of the integral of the brightest pixels in the panoramas, divided by the integral over all pixels.
-We further divide this integral by 2 to consider a single hemisphere and avoid overly bright dominant light sources.
-The light color is defined by , where is the normalized ambient color .
-We found that the bandwidth parameter did not advantage any specific method and therefore fixed it to , which we typically observe for an indoor light.
-4.3.2 Per-method details
-DiLightNet [71] DiLightNet requires a prompt describing the object, which we obtain by feeding the ground truth object composited on a black background to BLIP-2 [36] with the prompt “Putting aside the black background, this object is a”.
-To generate the results, depth and radiance hints are required.
-We supply DiLightNet with depth information from the 3D model of the object and render the radiance hints using the input environment map (sec.
-4.3.1) as outlined in [71].
-The rendered object is then composited with the background, along with the rendered shadow.
-Neural Gaffer [30] We provide the method with the environment map described in sec.
-4.3.1.
-The ground truth object is composited on a white background before feeding it to the network, as expected by the method.
-Since Neural Gaffer is trained at lower resolution (), its result is upsampled to then composited on the full resolution background.
-An erosion using a kernel is applied to the ground truth mask to mitigate boundary artifacts.
+We compare SpotLight with two types of baselines, that is methods that are conditioned on: the guiding shadow (like SpotLight); or an explicit lighting representation, through a environment map.
+We now describe per-method details.
+OpenImageDenoise [1] is also used to denoise the results of all methods to limit the impact of noise generated by diffusion models.
+5.3.1 Shadow-conditioned methods
+These methods, like ours, use the user-specified guiding shadow as input (obtained with EEVEE, see sec. 5.1).
+ZeroComp [78]+SDEdit [44] We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map.
+To blend the shadows, we perform a refinement step, inspired by SDEdit [44], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
 IC-Light [76] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning.
-We provide a rendering of the object lit by the ground truth environment map, giving a significant advantage to the method.
+We provide a rendering of the object lit by constant ambient lighting as input.
 No prompts are used.
+5.3.2 Light-conditioned methods
+Object relighting methods accept an explicit lighting representation as input in the form of a environment map, which we generate using the parametric light direction (see supp.).
+Unfortunately, these methods focus solely on object relighting—they do not create any cast shadows outside the object.
+To prevent this absence of shadows from severely hampering the realism of their results, we composite the guiding shadow onto the background image.
+DiLightNet [74] DiLightNet requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [35] with the prompt “Putting aside the black background, this object is a”.
+DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [74].
+The rendered object is then composited with the background.
+Neural Gaffer [29] The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method.
+Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background.
+We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
+Guiding shadow | IC-Light [76] | DiLightNet [74] | Neural Gaffer [29] | ZeroComp+SDEdit | Ours | Ground truth |
 5.4 Experimental results
@@
 sec.
-4.3) with two versions of SpotLight (c.f.
+5.3) with SpotLight (c.f.
 sec.
-4.2) on our “reference-based” dataset, where simulated ground truth is available (sec.
-4.1).
-Here, all methods are conditioned on the dominant light direction using their specific light parametrization (sec.
-4.3.1).
-It is worth noting that all three baselines (DiLightNet, IC-Light and Neural Gaffer) are given as input the ground truth object foreground and shadowed background, giving them a significant advantage.
+5.2) on our “reference-based” dataset, where the ground truth is available (sec.
+5.1).
+Here, all methods are conditioned on the ground truth dominant light direction using their specific light parametrization (sec.
+5.3).
 Fig.
@@
 We observe that SpotLight with the ZeroComp backbone outperforms the baselines on quantitative metrics and produces visually superior results.
-| Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
-| DiLightNet [71]∗ | 25.36 | 0.950 | 0.059 | 0.024 | 0.044 |
-| IC-Light [73]∗ | 28.59 | 0.965 | 0.043 | 0.018 | 0.034 |
-| Neural Gaffer [30]∗ | 29.64 | 0.963 | 0.037 | 0.016 | 0.038 |
-| SpotLight (RGBX) | 26.44 | 0.956 | 0.052 | 0.018 | 0.041 |
-| SpotLight (ZeroComp) | 30.68 | 0.973 | 0.033 | 0.012 | 0.031 |
-4.5 User study
-A user study was conducted to evaluate the realism of local lighting control.
+(a) Overall realism study
+(b) Lighting control study
+5.5 User studies
+Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines.
+In both cases, the Thurstone Case V Law of Comparative Judgement [61] is used to obtain a scaling -score for each method, where a higher scale indicates a higher human preference.
+Here, we give a high-level overview of those user studies.
+Refer to the supplementary material for additionnal details and studies.
+User study I: Overall realism
 Here, we use the “user-controlled” version of the dataset (see sec.
-4.1) to evaluate whether the realism of results obtained by conditioning on a specific lighting direction, even if it is not the dominant light direction.
-We evaluate the three methods with the best PSNR from tab.
-1, namely SpotLight (with the ZeroComp backbone), Neural Gaffer [30] and IC-Light [73].
+5.1) to evaluate the realism of results obtained by conditioning on a specific lighting direction, even if it is not aligned with the scene lighting conditions.
+We compare SpotLight against the three baselines with the best PSNR from tab.
+1, namely Neural Gaffer [29], IC-Light [76], and ZeroComp [78]+SDEdit [44].
 Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side.
-The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 40 random images are chosen, resulting in 120 pairs of images observed by each participant.
+The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant.
 We randomly shuffle the 120 pairs for each user, as well as the left-right order.
+Note that we also add three sentinel images to discard observers that did not understand the task.
 Examples are shown in fig.
-5.
-The user study was completed by observers.
-Following Giroux et al.
-[26], we used the Thurstone Case V Law of Comparative Judgement [60] to obtain a scaling -score for each method, where a higher scale indicates a higher human preference.
+3.
+User study I was completed by observers.
 The results, shown in fig.
-4 show that our method is the favorite.
-The 95% confidence interval error bars also show that the margin in performance is statistically significant, indicating that SpotLight achieves the highest realism.
-| Light direction | IC-Light [73] | DiLightNet [71] | Neural Gaffer [30] | Ours (RGBX) | Ours (ZeroComp) |
+4a, show that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
+User study II: lighting control Since the goal of our method is to give full artistic control over the object relighting, we design a second two-alternative forced choice user study to evaluate lighting controllability separately from the realism.
+The users were shown two videos side by side and were asked to “Select the video where the light direction has the most impact on the lighting of the object, by considering both the appearance of the object and its shadows.”.
+The videos show a virtual object relit by a rotating light source (see supp.).
+Here, we compared the two methods with the highest realism according to user study I: SpotLight and ZeroComp+SDEdit.
+Since this task aims to evaluate the local relighting within the object mask only, we replace the shadowed background with the output from our method.
+User study II was completed by observers.
+As seen in fig.
+4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy.
 5.6 Effect of parameter selection
 Our method employs parameters that control the degree of shadow guidance ( in eq.
-2) and the shadow blending ( in eq.
-1).
+3) and the shadow blending ( in eq.
+2).
 Tab.
-2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the simulated ground truth is available (c.f.
+2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the ground truth is available (c.f.
 sec.
-4.1).
+5.1).
 Although using no guidance () results in better quantitative performance, we observe qualitatively in fig.
 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control.
-Employing no shadow blending () results in similar metric values but makes shadows much less visible (fig.
-7).
+Employing no shadow blending () results in similar metric values but makes shadows much less visible (refer to the supp.
+for a qualitative comparison).
 Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
-| , (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.031 |
-| (no guidance) | 31.65 | 0.976 | 0.030 | 0.011 | 0.029 |
-| 28.60 | 0.966 | 0.044 | 0.015 | 0.036 | |
-| 30.78 | 0.974 | 0.033 | 0.012 | 0.030 | |
-| 29.32 | 0.970 | 0.039 | 0.014 | 0.034 |
+, (Ours) | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
+(no guidance) | 31.69 | 0.976 | 0.029 | 0.011 | 0.029 |
+28.68 | 0.966 | 0.043 | 0.015 | 0.036 | |
+30.81 | 0.974 | 0.032 | 0.012 | 0.029 | |
+29.24 | 0.969 | 0.039 | 0.014 | 0.034 |
 No guidance () | (ours) |
-| (no blending) | (ours) |
-5 Extensions
+6 Applications
 This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
-Shadow control on 2D object composition We demonstrate that SpotLight can be applied to 2D images by leveraging a ZeroComp [75] backbone trained without depth maps, using the normals [68] and albedo [33] estimators.
-Without access to a 3D model, we cannot rely on a rendering engine to cast shadows.
-Instead, we employ PixHt-Lab [57] to generate realistic soft shadows from a controllable point light coordinate.
-Here, a sample without shadow is provided for negative guidance , see sec.
-3.3.
-To estimate the required pixel height of the object cutout, we attribute each pixel -coordinate as the pixel height for demonstration purposes, more sophisticated methods such as [56] could be used.
-After rendering with the backbone, we enhance the target by adding the high-frequency details from the source image.
-This is achieved by subtracting a Gaussian-blurred version of the source image ( = 1) from itself (as in unsharp masking) and adding the resulting details to the target.
-Example 2D compositing results are shown in fig.
-8, which demonstrate realistic 2D object relighting from a user-controlled shadow.
-| Object | Background | Light-controllable composites | ||
-Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and pixel height in 2D.
+Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and, in the supp., pixel height in 2D [58].
 However, one can simply draw the desired shadow!
 Fig.
-9 and SpotLight: Shadow-Guided Object Relighting via Diffusion (first row) show that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
+6 shows that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
+Full image relighting As shown in fig.
+7, we can extend SpotLight to achieve full-scene relighting.
+Although our method is not explicitly designed for this task, we find our shadow guidance to be effective in plausibly relighting the background.
+First, we relight the object using SpotLight with the desired guiding shadow.
+Next, we apply a version of ZeroComp trained with an inverted shading mask, using only the object shading to relight the background.
+We repeat this process for different light directions and average the resulting predictions to obtain an average relit background.
+Finally, we compute a per-pixel highlight map by dividing each background prediction by the average and multiplying it by the original background.
+This process ensures that background shadows and highlights consistently align with the object’s shadow.
 7 Discussion
+Incorrect user-defined shadow | Correct shadow |
+Limitation: physically incorrect shadows Since SpotLight imposes no constraints on the input shadow, the user is free to specify a shadow direction that potentially contradicts the real lighting direction in the background image.
+In this case, SpotLight will still attempt to generate an image that looks as realistic as possible, given the shadow conditioning, as shown in fig.
+8.
 We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training.
 Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
-While the model can realistically estimate relighting from a shadow, generating a shadow map from scratch remains challenging.
+While the model can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some.
 A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights.

### v3

@@
 However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome.
-In this paper, we show that precise lighting control can be achieved for object relighting simply by providing a coarse shadow of the object.
+In this paper, we show that precise and controllable lighting can be achieved without any additional training, simply by supplying a coarse shadow hint for the object.
 Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image.
-Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training.
+Our method, SpotLight, is entirely training-free and leverages existing neural rendering approaches to achieve controllable relighting.
 We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting.
@@
 1 Introduction
-Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process.
-When one is given a photograph, however, no such control is possible.
-Relighting then becomes a hard task because of the complex interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture.
-The recent rise of powerful image generation models (e.g., [54]) has enabled exciting possibilities for image relighting [31, 76].
-This paper focuses on providing lighting control to recent diffusion image relighting frameworks, where one wishes to modify the appearance of a single object.
-This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image.
-In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
-We present SpotLight, a method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer.
-Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [74, 29]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer.
-Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it to relight the object accordingly.
-We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object as it moves around it.
-Fig.
-1 shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [19] according to the desired lighting direction.
+Object relighting—the process of inserting a virtual object in an image and lighting it appropriately—has traditionally been confined to 3D graphics pipelines, where lighting is explicitly modeled and rendered.
+In contrast, object relighting in real photographs remains a fundamentally ill-posed problem since it requires inferring the physical properties of the scene: its geometry, materials, and illumination are not observed, and their complex interactions—such as specular reflections, soft shadows, or material-dependent reflectance—must be inferred or approximated.
+Recent advances in image generative models, especially diffusion-based methods [rombach2022high], have opened up new possibilities for object relighting [zhang2025iclight, zhang2025zerocomp, liang2025diffusionrenderer].
+These models encode strong priors over the appearance of scenes under varied lighting, allowing implicit reasoning about illumination, material, and geometry.
+Yet, precise and localized relighting control remains limited, particularly in the context of virtual object insertion, where a single object must be relit.
+We focus on the problem of local lighting control for object compositing using generative models.
+Specifically, we aim to relight an inserted object such that the object conforms to a specified lighting direction and intensity around it, while blending it naturally with the background.
+Achieving such control is challenging: the relighting must align with physical cues like shadows and shading, yet remain flexible enough to work across varied materials, perspectives, and image contexts.
+Diffusion models offer powerful priors, but they are not inherently controllable without carefully designed guidance.
+We present SpotLight, a training-free method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer.
+Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [zeng2024dilightnet, jin2024neural_gaffer]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer.
+Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it.
+We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object to insert in the image.
+SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
+Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control.
+This mechanism is applied to recent diffusion models pre-trained to render images from input intrinsic maps (depth, normals, albedo, etc.) [zhang2025zerocomp, zeng2024rgb].
+SpotLight: Shadow-Guided Object Relighting via Diffusion shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [eisemann2011realtime] according to the desired lighting direction.
 From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting.
-To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
-Our core technical contribution lies in our shadow guidance approach, which refines the guiding shadow input while realistically relighting the virtual object.
-Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control.
-This builds upon existing object compositing networks [78, 75] that operate on intrinsic images along with diffusion models’ learned priors to render realistic outputs.
-We demonstrate that SpotLight’s guidance mechanism handles complex scenarios and maintains physical plausibility without requiring manual intervention nor any additional training.
 In summary, our contributions are as follows.
-First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring additional training.
+First, we introduce a framework that enhances pre-trained diffusion renderers with object relighting control without requiring additional training.
 We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models.
@@
 2 Related work
-Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [59, 27, 60, 54].
-These models offer control mechanisms for various modalities, such as text [50, 54], image content [55, 21], intrinsic images [78, 43, 75, 39], classifier guidance [16] and cross-modal data [12, 72].
-Notably, training-free guidance methods [44, 20, 3, 4, 26] have enabled image editing capabilities without specific training, forming the basis of our approach.
+Conditional image generation.
+Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [sohl2015deep, ho2020denoising, song2020denoising, rombach2022high].
+These models offer control mechanisms for various modalities, such as text [ramesh2022hierarchical, rombach2022high], image content [saharia2022palette, gal2022image], intrinsic images [zhang2025zerocomp, luo2024intrinsicdiffusion, zeng2024rgb, liang2025diffusionrenderer], classifier guidance [dhariwal2021diffusion] and cross-modal data [brooks2023instructpix2pix, ye2023ip].
+Notably, training-free guidance methods [meng2022sdedit, epstein2023diffusion, avrahami2022blended, avrahami2023blendedlatent, ho2021classifierfree] have enabled image editing capabilities without specific training, forming the basis of our approach.
 Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials.
-Previous methods for harmonization [53, 48, 34, 62] do not prioritize physically accurate shading of the objects.
-Recently, [49] utilized a single-view multi-illumination dataset [45] to enable direct control over the dominant lighting direction.
-Retinex theory has been leveraged [67, 11] for relighting indoor scenes.
-IC-Light [76] enforces consistency in appearance to relight portraits and various objects.
-LumiNet [68] and ScribbleLight [14] achieve image relighting through illumination transfer or user-defined scribbles, respectively.
-Some methods specialize in outdoor scenes, using geometric priors like depth [25] or normalized coordinates [31] or both [40].
+Previous methods for harmonization [reinhard2001color, pitie2005n, lalonde2007using, tsai2017deep] do not prioritize physically accurate shading of the objects.
+Recently, [poirier2024diffusion] utilized a single-view multi-illumination dataset [murmann2019multi] to enable direct control over the dominant lighting direction.
+Retinex theory has been leveraged [xing2024retinex, bhattad2024stylitgan] for relighting indoor scenes.
+IC-Light [zhang2025iclight] enforces consistency in appearance to relight portraits and various objects.
+LumiNet [xing2025luminet] and ScribbleLight [choi2025scribblelight] achieve image relighting through illumination transfer or user-defined scribbles, respectively.
+Some methods specialize in outdoor scenes, using geometric priors like depth [griffiths2022outcast] or normalized coordinates [kocsis2024lightit] or both [lin2025urbanir].
 These methods however cannot handle lighting on specific region or objects in the scene.
-Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13].
-Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [78, 75, 39], which can be leveraged to achieve zero-shot object composition.
+Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [bhattad2022cut] to harmonization of foreground-background albedo [careaga2023intrinsic].
+Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [zhang2025zerocomp, zeng2024rgb, liang2025diffusionrenderer], which can be leveraged to achieve zero-shot object composition.
 However, these methods have limited lighting control.
-ZeroComp [78] provides no lighting control, RGBX [75] can only provide lighting control through a text prompt, and DiffusionRenderer [39] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set.
-Other diffusion-based approaches, such as Neural Gaffer [29], DiLightNet [74] and IllumiNeRF [79] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images.
+ZeroComp [zhang2025zerocomp] provides no lighting control, RGBX [zeng2024rgb] can only provide lighting control through a text prompt, and DiffusionRenderer [liang2025diffusionrenderer] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set.
+Other diffusion-based approaches, such as Neural Gaffer [jin2024neural_gaffer], DiLightNet [zeng2024dilightnet] and IllumiNeRF [zhao2024illuminerf] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images.
+LightLab [magar2025lightlab] achieves fine-grained control over existing light sources, but does not handle object insertion.
+Careaga et al.
+[careagaRelighting] propose a sophisticated inverse rendering method for data generation, wheras SpotLight is training-free.
+IntrinsicEdit [lyu2025intrinsicedit] proposes a training-free method for lighting-based editing, but, unlike our approach, lacks explicit control over object-specific virtual lighting effects.
 We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
-Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 36, 80, 47, 38], though they generally lack controllability.
-Parametric models [23, 65, 15] or GAN inversion methods [64] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
-Shadow generation Generative models for rendering realistic shadows [77, 41, 28, 42], or harmonizing rendered shadows with the image [63] have been proposed.
-Other methods enable controllable shadows for 3D [56] and 2D objects [57, 58].
-Our approach can leverage existing shadow generation methods (e.g., [58], see supp.) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
-Image intrinsics Decomposing images into albedo and shading has long been studied [7, 33, 46, 66, 13].
-Numerous works have also focused on estimating scene geometry, for example, depth [51, 8, 73, 30, 69, 70] and normals [18, 6, 52, 5, 71].
-Recently, approaches such as [36, 80] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials.
-Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 30, 71, 2, 43].
+Explicit lighting estimation approaches infer HDR lighting from a single image [gardner2017sigasia, garon2019cvpr, li2020cvpr, zhu2022irisformer, phongthawee2024diffusionlight, liang2024photorealistic], though they generally lack controllability.
+Parametric models [gardner2019iccv, weber2022editable, dastjerdi2023everlight] or GAN inversion methods [wang2022stylelight] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
+Shadow generation.
+Generative models for rendering realistic shadows [zhang2019shadowgan, liu2020arshadowgan, hong2022shadow, liu2024shadow], or harmonizing rendered shadows with the image [valenca2023shadow] have been proposed.
+Other methods enable controllable shadows for 3D [sheng2021ssn] and 2D objects [sheng2022controllable, sheng2023pixht].
+Our approach can leverage existing shadow generation methods (e.g., [tasar2024controllable]) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
+Image intrinsics.
+Decomposing images into albedo and shading has long been studied [barrow1978recovering, kovacs2017shading, philip2019multi, wu2023measured, careaga2023intrinsic].
+Numerous works have also focused on estimating scene geometry, for example, depth [ranftl2020towards, bhat2023zoedepth, yin2023metric3d, ke2024repurposing, yang2024depth, depth_anything_v2] and normals [eftekhar2021omnidata, bae2021estimating, ranftl2021vision, bae2024rethinking, ye2024stablenormal].
+Recently, approaches such as [li2020cvpr, zhu2022irisformer] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials.
+Conversely, generative models have shown a powerful internal understanding of intrinsic properties [bhattad2024stylegan, du2023generative, ke2024repurposing, ye2024stablenormal, anonymous2024intrinsiccontrolnet, luo2024intrinsicdiffusion].
 We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
+| Guiding shadow | IC-Light [zhang2025iclight] | DiLightNet [zeng2024dilightnet] | Neural Gaffer [jin2024neural_gaffer] | ZeroComp+SDEdit | Ours | Ground truth |
 3 Background: diffusion renderers
-Recent work [75, 78] has shown that diffusion models can be adapted to generate images from intrinsic maps including materials (e.g., albedo, roughness, metallic), geometry (e.g., surface normals and depth), and shading.
-This capability was built using a ControlNet with ZeroComp [78] or full finetuning with RGBX [75].
+Recent work has shown that diffusion models can be adapted to generate photorealistic images from intrinsic maps including materials (e.g., albedo, roughness, metalness), geometry (e.g., surface normals and depth), and shading.
+This capability was built using a ControlNet in ZeroComp [zhang2025zerocomp] or full finetuning in RGBX [zeng2024rgb].
 We refer to this new class of conditional generation models as “diffusion renderers”.
-In these models, the denoising network is ran at each denoising step to obtain
-where represents the network -prediction, the noisy latents, and the conditioning intrinsic maps.
-It has been shown that these methods can realistically composite virtual objects into images in a zero-shot fashion, despite not being trained for this task.
-Provided a mask of the virtual object and its intrinsic maps (rendered with graphics shaders), the object can be trivially composited in most intrinsic maps of the background image (estimated with trained estimators).
-Since computing the virtual object shading requires complex lighting simulations, diffusion approaches estimate the shading by providing a partial shading map [78], or train inpainting models [75].
-However, unlike traditional rendering, these methods offer no control over the lighting conditions.
-In this work, we use a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, analogous to a spotlight moving around the object.
+Specifically, ZeroComp [zhang2025zerocomp], which we will use in sec.
+5, is trained to generate images from albedo, normals, depth, and partially-masked shading.
+It can be used for compositing in a zero-shot manner by alpha-compositing the intrinsics of the background and the object, and by masking out a region on and around the object in the shading map.
+The network then generates the full image with realistic shading.
+The intrinsics of the object are rendered using simple graphics shaders, while those of the background are estimated using pre-trained networks (e.g., [bhat2023zoedepth] for depth, see sec.
+5.2).
+In the case of RGBX [zeng2024rgb], their XRGB network achieves object insertion by accepting as input alpha-composited albedo, normals, and a masked image.
+They then train an inpainting version of their network to generate the final image.
+Unlike traditional rendering, these methods offer no control over the lighting conditions.
+In this work, we use such a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, without any additional training.
 We now describe our approach in detail.
 4 SpotLight
-We employ a user-provided, guiding shadow to steer the pre-trained diffusion process, thereby enforcing matching illumination on the inserted object.
-Our key insight is that even an approximate shadow encapsulates important lighting information.
-By integrating this cue early in the diffusion process, we simultaneously refine both the composite appearance and its shadow, yielding a natural blend with the scene.
+We employ a user-provided, guiding shadow to steer a pre-trained diffusion renderer (c.f.
+sec.
+3) to follow the desired lighting direction.
+Our key insight is that even an approximate shadow encapsulates enough important lighting information and, by integrating this cue in the denoising process, we can simultaneously refine the composite appearance and its shadow, yielding a natural blend with the scene.
 4.1 Approach overview
-Our approach requires two inputs: an object mask and a user-specified guiding shadow .
-This shadow, computed via fast rasterization or specified through hand-drawn scribbles, is used to guide the latent diffusion process.
-Without retraining the pre-trained renderer , we steer its generative process by incorporating the guiding shadow into the latent space.
-4.2 Blending shadows
-To fuse the guiding shadow with the latent representation, we update the noisy latent code at each diffusion timestep, similar to [4]. In our framework, the updated latent
-serves as a soft constraint that steers the process toward a shadow-consistent output.
-Here, is the composite of the object albedo, the target shadow, and the background, and controls the strength of the guiding shadow.
-The latent shadow mask is obtained by downsampling the target shadow to the latent resolution (see supp.
+An overview of SpotLight is illustrated in fig.
+2.
+In addition to the intrinsic maps for the backbone diffusion renderer, our approach requires two inputs: an object mask and a user-specified guiding shadow which guides the latent diffusion process.
+The guiding shadow can be obtained in several ways—in this paper, we experiment with fast rasterization [eisemann2011realtime], a trained shadow generation model [tasar2024controllable], and hand-drawn scribbles.
+Without retraining the diffusion renderer, we steer its generative process by incorporating the guiding shadow into the latent space.
+4.2 Blending shadows with the background
+Provided the guiding shadow , we first need to ensure that the diffusion model will properly blend the shadow with the background.
+We take inspiration from Blended Latent Diffusion [avrahami2023blendedlatent] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadow roughly composited over the image, .
+In practice, we obtain by compositing the object albedo and the guiding shadow on the background.
+The update made at each timestep to incorporate the desired shadow latents is
+where is the shadow latent weight and is the shadow mask, bilinearly downsampled to the latent resolution (), with the edge opacity increased and dilated to maintain the desired shadow softness (see supp.
 for details).
-4.3 Shading the object via dual-branch guidance
-A key element of our approach is to amplify the impact of the guiding shadow on the object shading through classifier-free guidance (CFG) [26].
-We process two parallel diffusion branches: a positive branch that incorporates the desired shadow direction, and a negative branch that suppresses an undesired shadow direction (e.g., via an opposing light direction).
-Fig.
-2 shows our dual-branch guidance mechanism.
-Classifier-free guidance then combines these branches:
-where sets the guidance scale and is the object mask downsampled to the latent resolution.
-This dual-branch strategy reinforces the intended lighting on the object while naturally blending it with the background.
+Noise is added to the encoded composite using the DDIM [song2020denoising] noise scheduler.
+Different from [avrahami2023blendedlatent], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
+4.3 Enhancing lighting control
+Simply running a diffusion renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec.
+5.6).
+We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by .
+Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample.
+From the outputs of the diffusion renderer (typically, -predictions [salimans2022progressive], i.e., linear combination of noise and estimated image), we then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
+where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After eq. 2, we run a regular DDIM sampling step to obtain the latents .
 4.4 Final image synthesis
-After the guided diffusion process, the resulting latent is decoded with the VAE decoder.
-The background preservation strategy of [78] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
+After the guided diffusion process, the resulting latents are decoded with the VAE decoder.
+The background preservation strategy of [zhang2025zerocomp] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
 5 Evaluation on 3D object compositing
 Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene.
-To address this, we build over the evaluation setup of [78] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas.
+To address this, we build over the evaluation setup of [zhang2025zerocomp] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas.
 We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render.
@@
 5.1 Evaluation dataset
-We create datasets for the two scenarios using the Blender scene information graciously provided by the authors of [78].
-We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [70].
-We also improve the ground truth intrinsic maps by providing correct antialiased versions.
-“Reference-based” scenario To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [78] to ensure lighting fidelity around the inserted object.
-We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [19].
+We create datasets for the two scenarios using the Blender scenes graciously provided by the authors of [zhang2025zerocomp].
+We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [depth_anything_v2].
+We also improve the ground truth intrinsic maps with antialiased versions.
+“Reference-based” scenario.
+To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [zhang2025zerocomp] to ensure lighting fidelity around the inserted object.
+We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [eisemann2011realtime].
 A total of 210 images featuring various objects are rendered, with example samples shown in fig.
@@
 5.4).
-“User-controlled” scenario This dataset simulates a scenario where a user specifies a desired light position.
+“User-controlled” scenario.
+This dataset simulates a scenario where a user specifies a desired light position.
 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine.
 The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images.
-This dataset is utilized in the user study (sec.
+This dataset is used in the user study (sec.
 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.
@@
 As mentioned in sec.
-3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [78] or RGBX [75].
+3, our proposed SpotLight readily applies to existing diffusion-based neural renderers, such as ZeroComp [zhang2025zerocomp] or RGBX [zeng2024rgb].
 Here, we provide details on ZeroComp, see supp.
 for RGBX.
-We use the checkpoint trained on the Openrooms dataset [37] provided by the authors.
+We use the variant trained on Openrooms [li2020openrooms] provided by the authors.
 The masked shading map is obtained by dividing the background image by the albedo, and by masking out the shading on both the object and shadow regions.
 Given its superior results, we adopt this method as our default model for 3D object compositing.
-To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [78], namely StableNormal [71] for normals, IID [32] for albedo, and ZoeDepth [8] for depth.
+To extract intrinsics from real images, we use the same conditioned intrinsic estimators as in [zhang2025zerocomp], namely StableNormal [ye2024stablenormal] for normals, IID [kocsis2023iid] for albedo, and ZoeDepth [bhat2023zoedepth] for depth.
 | Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
 | Light-conditioned methods | |||||
-| DiLightNet [74] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
-| Neural Gaffer [29] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
+| DiLightNet [zeng2024dilightnet] | 24.67 | 0.948 | 0.064 | 0.022 | 0.042 |
+| Neural Gaffer [jin2024neural_gaffer] | 28.44 | 0.963 | 0.042 | 0.015 | 0.038 |
 | Shadow-conditioned methods | |||||
-| IC-Light [76] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
-| ZeroComp+SDEdit [44] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
+| IC-Light [zhang2025iclight] | 26.87 | 0.959 | 0.054 | 0.019 | 0.040 |
+| ZeroComp+SDEdit [meng2022sdedit] | 26.00 | 0.938 | 0.053 | 0.025 | 0.079 |
 | SpotLight | 30.68 | 0.974 | 0.033 | 0.012 | 0.030 |
@@
 We now describe per-method details.
-OpenImageDenoise [1] is also used to denoise the results of all methods to limit the impact of noise generated by diffusion models.
+OpenImageDenoise [OpenImageDenoise] is also used to denoise the results of all methods to limit the impact of rendering noise present in the synthetic training data of some models.
 5.3.1 Shadow-conditioned methods
 These methods, like ours, use the user-specified guiding shadow as input (obtained with EEVEE, see sec. 5.1).
-ZeroComp [78]+SDEdit [44] We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map.
-To blend the shadows, we perform a refinement step, inspired by SDEdit [44], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
-IC-Light [76] We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning.
-We provide a rendering of the object lit by constant ambient lighting as input.
+ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit].
+We create a strong baseline by first running ZeroComp with the shading of the object masked out, and composite the guiding shadow over the background shading intrinsic map.
+To blend the shadows, we perform a refinement step, inspired by SDEdit [meng2022sdedit], by noising this prediction for 50% of the timesteps, and re-running ZeroComp for the remaining timesteps, this time with the shading of the whole image masked out.
+IC-Light [zhang2025iclight].
+We evaluate the background-conditioned model of IC-Light, where the background image with the shadow is used for conditioning, with the object lit by a constant ambient lighting as input.
 No prompts are used.
@@
 To prevent this absence of shadows from severely hampering the realism of their results, we composite the guiding shadow onto the background image.
-DiLightNet [74] DiLightNet requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [35] with the prompt “Putting aside the black background, this object is a”.
-DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [74].
+DiLightNet [zeng2024dilightnet] requires a prompt describing the object, which we obtain by feeding the object lit by constant ambient lighting and composited on a black background to BLIP-2 [li2023blip2] with the prompt “Putting aside the black background, this object is a”.
+DiLightNet also requires depth and radiance hints: we obtain depth from the 3D model of the object and render the radiance hints using the input environment map as outlined in [zeng2024dilightnet].
 The rendered object is then composited with the background.
-Neural Gaffer [29] The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method.
+Neural Gaffer [jin2024neural_gaffer].
+The object lit by constant ambient lighting is composited on a white background before feeding it to the network, as expected by the method.
 Since Neural Gaffer is trained at a lower resolution (), its result is upsampled to then composited on the full resolution background.
 We noticed Neural Gaffer tends to generate noticeable white boundary artifacts—we remove them through an erosion operation to avoid overly penalizing it.
-| Guiding shadow | IC-Light [76] | DiLightNet [74] | Neural Gaffer [29] | ZeroComp+SDEdit | Ours | Ground truth |
 5.4 Experimental results
@@
 Two user studies were conducted to assess the perceptual realism and controllability of SpotLight compared to baselines.
-In both cases, the Thurstone Case V Law of Comparative Judgement [61] is used to obtain a scaling -score for each method, where a higher scale indicates a higher human preference.
-Here, we give a high-level overview of those user studies.
-Refer to the supplementary material for additionnal details and studies.
+In both cases, the Thurstone Case V Law of Comparative Judgement [thurstone2017] is used to obtain a -score for each method, where a higher value indicates higher human preference.
+Here, we give an overview of those user studies.
+Refer to the supp.
+for additional details and studies.
 User study I: Overall realism
@@
 We compare SpotLight against the three baselines with the best PSNR from tab.
-1, namely Neural Gaffer [29], IC-Light [76], and ZeroComp [78]+SDEdit [44].
+1, namely Neural Gaffer [jin2024neural_gaffer], IC-Light [zhang2025iclight], and ZeroComp [zhang2025zerocomp]+SDEdit [meng2022sdedit].
 Since no ground truth is available, we design a two-alternative forced choice user study, where results obtained by two methods are shown side by side.
 The participants are asked to “Click on the image that looks the most realistic, by considering both the appearance of the object and its shadows.” For each pair of methods, a set of 20 random images are chosen (we ensure images where shadows are clearly visible are selected), resulting in 120 pairs of images observed by each participant.
-We randomly shuffle the 120 pairs for each user, as well as the left-right order.
-Note that we also add three sentinel images to discard observers that did not understand the task.
+We randomly shuffle the 120 pairs for each user, and the left-right order.
+Three sentinel images were added to discard observers that did not understand the task.
 Examples are shown in fig.
@@
 User study I was completed by observers.
-The results, shown in fig.
-4a, show that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
+Fig.
+4a shows that our method is the favorite, indicating that SpotLight achieves the highest realism, with a statistically significant difference to Neural Gaffer and IC-Light.
 User study II: lighting control Since the goal of our method is to give full artistic control over the object relighting, we design a second two-alternative forced choice user study to evaluate lighting controllability separately from the realism.
@@
 4b, SpotLight achieves a statistically significant improvement in performance over ZeroComp+SDEdit, despite the low number of participants, which we attribute to the use of our local guidance strategy.
+Please see the supp.
+for more user study results and experimental details.
 5.6 Effect of parameter selection
-Our method employs parameters that control the degree of shadow guidance ( in eq.
-3) and the shadow blending ( in eq.
-2).
+Our method uses two tunable parameters: the local guidance scale ( in eq.
+2) and the shadow blending weight ( in eq.
+1).
 Tab.
-2 shows the impact of parameter selection on quantitative metrics, obtained on the “reference-based” dataset where the ground truth is available (c.f.
+2 reports metrics obtained on the “reference-based” dataset where the ground truth is available (c.f.
 sec.
@@
 5 that this significantly reduces the visibility of generated shading on the inserted object, thereby reducing the impact of the desired local lighting control.
-Employing no shadow blending () results in similar metric values but makes shadows much less visible (refer to the supp.
-for a qualitative comparison).
+Employing no shadow blending () results in similar metric values but makes shadows much less visible.
+Please refer to the supp.
+for a qualitative comparison for the , and the shadow softness (light radius) parameters.
 | Method | PSNR↑ | SSIM↑ | RMSE↓ | MAE↓ | LPIPS↓ |
@@
 This section highlights additional capabilities derived from our framework to further demonstrate its versatility.
-Scribbles as shadows Shadows can be generated in multiple ways: we have demonstrated the use of shadow mapping in 3D and, in the supp., pixel height in 2D [58].
-However, one can simply draw the desired shadow!
+Scribbles as shadows.
+Instead of relying on shadow mapping in 3D, one can simply draw the desired shadow!
 Fig.
-6 shows that SpotLight is able to realistically relight the object and refine its shadow even when it is a user-drawn scribble.
-Full image relighting As shown in fig.
-7, we can extend SpotLight to achieve full-scene relighting.
-Although our method is not explicitly designed for this task, we find our shadow guidance to be effective in plausibly relighting the background.
-First, we relight the object using SpotLight with the desired guiding shadow.
-Next, we apply a version of ZeroComp trained with an inverted shading mask, using only the object shading to relight the background.
-We repeat this process for different light directions and average the resulting predictions to obtain an average relit background.
-Finally, we compute a per-pixel highlight map by dividing each background prediction by the average and multiplying it by the original background.
-This process ensures that background shadows and highlights consistently align with the object’s shadow.
+6 shows that SpotLight can realistically relight the object and refine its shadow even when it is a user-drawn scribble.
+2D objects.
+SpotLight can be applied to real 2D objects from existing photographs by estimating their intrinsics, as in sec.
+5.2, and generating desired shadows with shadow generation methods such as [tasar2024controllable], as illustrated in fig.
+7.
+| Input object | Relit object on target background | ||
+Reflective materials.
+When using a ZeroComp backbone pre-trained on the InteriorVerse dataset [zhu2022learning], SpotLight naturally generalizes to metallic objects and generates controllable specular reflections on the object, see fig.
+8.
+Additional relighting results.
+SpotLight also shows other capabilities: relighting for more diverse objects and outdoor scenes, simulating multiple light sources by combining outputs with different guiding shadows, and full-image relighting.
+Due to space constraints, the implementation details and results are provided in the supp.
 7 Discussion
-| Incorrect user-defined shadow | Correct shadow |
-Limitation: physically incorrect shadows Since SpotLight imposes no constraints on the input shadow, the user is free to specify a shadow direction that potentially contradicts the real lighting direction in the background image.
-In this case, SpotLight will still attempt to generate an image that looks as realistic as possible, given the shadow conditioning, as shown in fig.
-8.
-We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without requiring any additional training.
+| Incorrect user-defined shadow | Correct shadow | ||
+Limitation: physically incorrect shadows.
+Since SpotLight imposes no constraints on the input shadow, users may specify a shadow direction that potentially contradicts the real lighting in the background.
+In this case, SpotLight still attempts to render a visually plausible composition consistent with the given shadow, as shown in fig.
+9.
+We present SpotLight, a method for realistically controlling the local lighting of an object through a coarse shadow, compatible with diffusion-based intrinsic image renderers without any additional training.
 Our results demonstrate that fine-grained control over local lighting can be achieved while attaining realistic compositions.
-While the model can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some.
+While our method can realistically generate relighting results from a shadow and despite showing that this can be done in multiple ways (shadow mapping, 2D generation, hand-drawn scribbles), the process of generating a shadow from scratch could prove challenging to some.
 A promising future direction is to enable end-to-end shadow control using parametric light models, such as point or sphere lights.
 Additionally, similar approaches could be explored to enable global lighting adjustments for entire images.
-Acknowledgements
-This research was supported by NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix.
+Acknowledgements.
+This research was supported by FRQNT scholarship 328810, NSERC grants RGPIN 2020-04799 and ALLRP 586543-23, Mitacs and Depix.
 Computing resources were provided by the Digital Research Alliance of Canada.
-The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all members of the lab for discussions and proofreading help.
+The authors thank Zheng Zeng, Yannick Hold-Geoffroy and Justine Giroux for their help as well as all lab members for discussions and proofreading help.